OpenStudy (anonymous):

What is the geometric mean of 6 and 13?

OpenStudy (the_fizicx99):

Geometric mean = (a1 × a2 ... an)1/n (6 * 13)1/2 78 * 1/2 = 39

random231 (random231):

u r wrong

OpenStudy (the_fizicx99):

I did square it but got a really long decimal, so I thought it was wrong. √78

OpenStudy (the_fizicx99):

You could use an exponent, (78)^1/2

OpenStudy (whpalmer4):

it's \[(a_1*a_2*...*a_n)^{1/n}\] The first attempt (=39) is incorrect. If you think about it, the geometric mean must be somewhere between the smallest and largest numbers.

OpenStudy (the_fizicx99):

...I know its incorrect. I corrected it.

OpenStudy (whpalmer4):

Yes, you did. I'm trying to help anyone reading this have an intuitive grasp of what the answer should (or shouldn't) be, to avoid any such future mistakes :-)

OpenStudy (the_fizicx99):

Lol so in a way, I showed what no to do (>.<) not exactly satisfying, but ok.